# LEP3: A High-Luminosity $e^{+}e^{-}$ Higgs & Electroweak Factory in the LHC Tunnel

### v2

@@ 3 Technical Considerations
Power Consumption: We estimate that the total power consumption at the highest LEP3 energy will be about 250 $MW$, and about 185 $MW$ at the Z pole.
@@ 4.1 Higgs Boson Physics
Table 4: The precision for Higgs boson couplings measurements. The errors for LEP3 have been increased with respect to the recent FCC-ee submission  [9], by a factor $\sqrt{22/4.2}$. In red are HL-LHC values that are unlikely to be improved by $e^{+}e^{-}$ Higgs boson factories.
@@ 4.2 Electroweak and Flavour Physics
With $1.7\times 10^{12}$ Z boson decays recorded, we expect $2.1\times 10^{11}B^{0}/\bar{B}^{0},2.1\times 10^{11}B^{+}/B^{-},5.1\times 10^{10}B_{s}^{0},1.3\times 10^{9}B_{c}$ decays, as well as $1.7\times 10^{12}$ charm quark pairs and $6\times 10^{10}\tau^{+}\tau^{-}$ pairs. This enables precision and extensive flavour physics programme at LEP3, largely complementary to the HL-LHC and Belle II programs. In particular, precision measurements of fundamental parameters of tau leptons (lifetime, mass, leptonic branching fraction), rare $B_{s}^{0}$ and $B_{c}$ decays and b baryon decays with missing energy, and charm physics will yield very rich and often unique results.
@@ 5 The Longer-Term Future
A major objective in particle physics is always to operate an accelerator that allows a leap of an order of magnitude in the constituent centre-of-mass (CoM) energy with respect to the previous one. Following on from the LHC, which operates at constituent $\sqrt{s}$ $\sim$ 1 $TeV$, the two possibilities to attain $\sqrt{s}\sim$ 10 $TeV$ directly are the FCC-hh operating at $\sqrt{s}$ $\sim$ 100 $TeV$ or a muon collider operating at $\sqrt{s}\geq$ 10 $TeV$.
@@
Whatever is the accelerator covering this gap, we expect that it would have to be built within the financial resources currently available to CERN and its particle physics user community. Any proposed project should address the recommendation (c) from the 2019 ESPP, which we interpret as requiring the production of a large sample of Higgs bosons for a more powerful analysis than is currently possible. It should also be able to explore the 10 $TeV$ scale indirectly via a suite of high precision electroweak measurements. For these reasons, we think that the LEP3 accelerator proposed here would provide an ideal back-up in the event of technical or financial issues for the preferred FCC project.
@@ 6 Summary
We support FCC as the preferred option for CERN’s future, but recall that the guidance for the 2026 ESPP strategy review requests proposals for alternatives to this preferred option. Linear e+e-, circular muon and LHeC colliders are among the alternatives that have been proposed. We propose that another option, an e+e-collider in the LHC tunnel, referred to here as LEP3, also be considered as an alternative. We have described its capabilities as an e+e- Higgs and electroweak factory. Compared to the linear e+e- colliders proposed, LEP3 with two effective experiments would make measurements with similar precision in the ZH mode and with superior precision at the Z and WW energies, all at much lower cost. No showstoppers have yet been identified, and we consider this proposal to be sufficiently interesting to deserve further study. That said, we have identified important areas that would require deeper investigation before CERN could commit to LEP3.

### v3

@@ 3 Technical Considerations
Power Consumption: We estimate that the total power consumption at the highest LEP3 energy will be about 250 $MW$, and about 185 $MW$ at the Z pole. It is likely that low temperature (LTS) or high temperature (HTS) superconducting quadrupole and sextupole magnets will be needed to minimise power consumption.
@@ 4.1 Higgs Boson Physics
Table 4: The precision for Higgs boson couplings measurements. The errors for LEP3 have been increased with respect to the recent FCC-$e$ submission  [8], by a factor $\sqrt{22/4.2}$. In red are HL-LHC values that are unlikely to be improved by $e^{+}e^{-}$ Higgs boson factories.
@@ 4.2 Electroweak and Flavour Physics
We expect to record $1.7\times 10^{12}$ visible Z boson decays, including $2.1\times 10^{11}B^{0}/\bar{B}^{0},2.1\times 10^{11}B^{+}/B^{-},5.1\times 10^{10}B_{s}^{0},1.3\times 10^{9}B_{c}$ decays, as well as $1.7\times 10^{12}$ charm quark pairs and $6\times 10^{10}\tau^{+}\tau^{-}$ pairs. This will enable an extensive high-precision flavour physics programme at LEP3, largely complementary to the HL-LHC and Belle II programmes. In particular, precision measurements of fundamental parameters of tau leptons (lifetime, mass, leptonic branching fraction), rare $B_{s}^{0}$ and $B_{c}$ decays and b baryon decays with missing energy, and charm physics will yield very rich and often unique results.
@@ 5 The Longer-Term Future
A major objective in particle physics is always to operate an accelerator that allows a leap of an order of magnitude in the constituent centre-of-mass (CoM) energy with respect to the previous one. Following on from the LHC, which operates at constituent $\sqrt{s}$ $\sim$ 1 $TeV$, the two possibilities to attain $\sqrt{s}\sim$ 10 $TeV$ directly are the FCC-hh operating at $\sqrt{s}$ $\sim$ 100 $TeV$ or a muon collider operating at $\sqrt{s}\sim$ 10 $TeV$.
@@
Whatever is the accelerator covering this gap, we expect that it would have to be built within the financial resources currently available to CERN and its particle physics user community. Any proposed project should address the recommendation (c) from the 2020 ESPP, which we interpret as requiring the production of a large sample of Higgs bosons for a more powerful analysis than is currently possible. It should also be able to explore the 10 $TeV$ scale indirectly via a suite of high precision electroweak measurements. For these reasons, we think that the LEP3 accelerator proposed here would provide an ideal back-up in the event of technical or financial issues for the preferred FCC project.
@@ 6 Summary
We support FCC as the preferred option for CERN’s future, but recall that the guidance for the 2026 ESPP strategy review requests proposals for alternatives to this preferred option. Linear e+e-, circular muon and LHeC colliders are among the alternatives that have been proposed. We propose that another option, an e+e-collider in the LHC tunnel, referred to here as LEP3, also be considered as an alternative, possibly as the primary alternative. We have described its capabilities as an e+e- Higgs and electroweak factory. Compared to the linear e+e- colliders proposed, LEP3 with two effective experiments would make measurements with similar precision in the ZH mode and with superior precision at the Z and WW energies, all at much lower cost. No showstoppers have yet been identified, and we consider this proposal to be sufficiently interesting to deserve further study and investment. That said, we have identified important areas that would require deeper investigation before CERN could commit to LEP3.